# Gradient Methods with Online Scaling

### v1

@@ 1. A1:
    $f(x)$ is $L$-smooth. $|f(x)-f(y)-\langle\nabla f(y),x-y\rangle|\leq\tfrac{L}{2}\|x-y\|^{2}$.
@@ 2. A2:
    $f(x)$ is $\mu$-strongly convex. $f(x)-f(y)-\langle\nabla f(y),x-y\rangle\geq\frac{\mu}{2}\|x-y\|^{2}$.
@@ 7 Practical considerations
In this section, we consider practical aspects of implementing OSGM in practice.
7.1 Gradient evaluations for $r_{x}$ and $h_{x}$
A first look at Algorithm 2 and Algorithm 4 suggests an additional gradient evaluation at every iteration. Both $\nabla r_{x}$ and $\nabla h_{x}$ need to evaluate two gradients in every iteration. However, Using the update rule $x^{k+1}=x^{k}-P_{k}\nabla f(x^{k})$, the gradient $\nabla r_{x^{k}}(P)$ in Algorithm 2 can be expressed as
@@ 7.1 Gradient evaluations for $r_{x}$ and $h_{x}$
If $x^{k+1}=x^{k+1/2}$, then $\nabla f(x^{k+1/2})$ can be reused in the next iteration. This is always true for $r_{x}$ when the monotone oracle is not used ($\mathcal{M}=\varnothing$), and the total number of gradient evaluations in OSGM-R is the same as that of gradient descent but simply requires an additional cache to store $\nabla f(x^{k+1})$. Regarding $\nabla g_{x}$, its implementation needs a Hessian-gradient product, which can be efficiently computed in practice.
@@ 7.2 Efficient scaling matrix updates
The subset $\mathcal{P}$ can be chosen to have a simple structure, such as diagonal matrices or sparse matrices with some prespecified sparsity pattern. Then the scaling matrix $P_{k+1}$ can be efficiently updated in the cost of $\mathcal{O}(\texttt{supp}(\mathcal{P}))$ since it suffices to compute the non-zero entries of the sparsity pattern in $\mathcal{P}$. For example, if $\mathcal{P}$ is the set of diagonal matrices, then (32) simplifies to $\nabla r_{x^{k}}(P)=-\tfrac{\nabla f(x^{k+1/2})\odot\nabla f(x^{k})}{f(x^{k})-f(x^{\star})}$, where $\odot$ denotes the element-wise product. A simpler structure in $\mathcal{P}$ provides more efficient scaling matrix updates. However, more freedom in $\mathcal{P}$ may provide smaller $\theta^{\star}_{K}$, enhancing the convergence of OSGM.
@@ 7.4 Choice of online learning algorithm $\mathcal{A}$
Our convergence analyses show that a good online learning algorithm $\mathcal{A}$ benefits convergence of OSGM. For simplicity, the simplest possible online learning algorithms are adopted in the theoretical analysis. However, OSGM is compatible with more advanced online learning algorithms such as AdaGrad. Using advanced online algorithms often greatly improves the robustness and practical performance of OSGM. In particular, our results on the hypergradient surrogate loss provide new insights into improving hypergradient descent heuristics.
@@ Appendix D Proof of results in Section 5
$\|\nabla f(x^{k+1})\|=\|\nabla f(\mathcal{M}(x^{k+1/2}))\|\leq\|\nabla f(x^{k+1/2})\|=g_{x^{k}}(P_{k})\|\nabla f(x^{k})\|.$
@@ Appendix E Proof of results in Section 6
Since $f(x^{k+1})=f(\mathcal{M}(x^{k+1/2}))\leq\min\{f(x^{k}),f(x^{k+1/2})\}$, we have

### v2

@@ 1. A1:
    $f(x)$ is $L$-smooth. $|f(x)-f(y)-\langle\nabla f(y),x-y\rangle|\leq\tfrac{L}{2}\|x-y\|^{2}$
@@ 2. A2:
    $f(x)$ is $\mu$-strongly convex. $f(x)-f(y)-\langle\nabla f(y),x-y\rangle\geq\frac{\mu}{2}\|x-y\|^{2}$
@@ 4.2 Online learning algorithm
Remark 1.
The relation (11) suggests additional complexity from a rank-one update with an orthogonal projection. But as we will discuss in Section 7, we can choose $\mathcal{P}$ to have arbitrary sparsity (e.g., diagonal), and it is only necessary to update the nonzero elements. Moreover, the orthogonal projection is often easy to compute since we do not require $P_{k}$ to be positive semidefinite or symmetric.
@@ 7 Practical considerations
This section considers practical aspects of implementing OSGM.
7.1 Gradient evaluations for $r_{x}$ and $h_{x}$ using simple comparison oracle
A first look at Algorithm 2 and Algorithm 4 suggests an additional gradient evaluation at every iteration. Both $\nabla r_{x}$ and $\nabla h_{x}$ need to evaluate two gradients in every iteration. However, with $x^{k+1/2}:=x^{k}-P_{k}\nabla f(x^{k})$ and simple comparison (Section 2.3) as the monotone oracle, the gradient $\nabla r_{x^{k}}(P)$ in Algorithm 2 can be expressed as
@@ 7.1 Gradient evaluations for $r_{x}$ and $h_{x}$ using simple comparison oracle
If $x^{k+1}=x^{k+1/2}$, then $\nabla f(x^{k+1/2})$ can be reused in the next iteration; if $x^{k+1}=x^{k}$, then $\nabla f(x^{k})$ can be reused. Therefore, simple comparison oracle ensures that the total number of gradient evaluations in OSGM-R is the same as that of gradient descent but simply requires an additional cache to store $\nabla f(x^{k+1})$. Regarding $\nabla g_{x}$, its implementation needs a Hessian-gradient product, which can be efficiently computed in practice.
@@ 7.2 Efficient scaling matrix updates
The subset $\mathcal{P}$ can be chosen to have a simple structure, such as diagonal matrices or sparse matrices with some prespecified sparsity pattern. Then the scaling matrix $P_{k+1}$ can be efficiently updated in the cost of $\mathcal{O}(\texttt{supp}(\mathcal{P}))$ since it suffices to compute the non-zero entries of the sparsity pattern in $\mathcal{P}$. For example, if $\mathcal{P}$ is the set of diagonal matrices, then (32) simplifies to $\nabla r_{x^{k}}(P)=-\tfrac{\nabla f(x^{k}-P_{k}\nabla f(x^{k}))\odot\nabla f(x^{k})}{f(x^{k})-f(x^{\star})}$, where $\odot$ denotes the element-wise product. A simpler structure in $\mathcal{P}$ provides more efficient scaling matrix updates. However, more freedom in $\mathcal{P}$ may provide smaller $\theta^{\star}_{K}$, enhancing the convergence of OSGM.
@@ 7.4 Choice of online learning algorithm $\mathcal{A}$
Our convergence analyses show that a good online learning algorithm $\mathcal{A}$ benefits convergence of OSGM. For simplicity, the simplest possible online learning algorithms are adopted in the theoretical analysis. However, OSGM is compatible with more advanced online learning algorithms such as AdaGrad. Using advanced online algorithms often greatly improves the robustness and practical performance of OSGM. In particular, our results on the hypergradient surrogate loss provide new insights into improving the hypergradient descent heuristics.
@@ Appendix D Proof of results in Section 5
$\|\nabla f(x^{k+1})\|=\|\nabla f(\mathcal{M}(x^{k}))\|\leq\|\nabla f(x^{k}-P_{k}\nabla f(x^{k}))\|=g_{x^{k}}(P_{k})\|\nabla f(x^{k})\|.$
@@ Appendix E Proof of results in Section 6
For convenience we denote $x^{k+1/2}:=x^{k}-P_{k}\nabla f(x^{k})$. By definition of the monotone oracle, we always have
$f(x^{k+1})=f(\mathcal{M}(x^{k}))\leq\min\{f(x^{k}),f(x^{k+1/2})\}.$
@@
Since $f(x^{k+1})=f(\mathcal{M}(x^{k}))\leq\min\{f(x^{k}),f(x^{k+1/2})\}$, we have